OpenStudy (jennyrlz):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (anonymous):

need help?

OpenStudy (jennyrlz):

yes c:

OpenStudy (anonymous):

one thing to know is that the vertex of a parabola is half-way between the focus and the directrix. This means the x-value of the vertex (the point where the parabola turns) of the parabola is the same as the focus and the directrix (a straight line), also has a point on it with the same x-value. Since the focus is at (0, 2) and directrix is y = -2, then the point where both of these have the same x-value will be at (0, 2) for the focus and (0, -2) for the directrix. The vertex will also have the same x-value so it will be (0, y). I don't know what the y-value of the vertex is, but I know that it's y-value is half-way between the y-value of the focus, and the y-value of the directrix at x = 0. Directrix y-value is -2 at x = 0 and for the focus it's 2 at x = 0. Halfway between y = -2 and y = 2 is y = 0. So the vertex of the parabola occurs at (0, 0). So that's x^2 = 4ay = 4(2)y = 8y. Rearranging gives: y = 1/8*x^2 2

OpenStudy (jennyrlz):

ok so you got 2 by dividing the given four by 2 correcy?

OpenStudy (anonymous):

yes :)

OpenStudy (jennyrlz):

xD

OpenStudy (jennyrlz):

um... do you hapen to know the formula for a parabola? i cant remember it >.<

OpenStudy (anonymous):

vertex formula???

OpenStudy (jennyrlz):

also was that it, the answer 0.o if so it doesnt match any of my answer choices T^T

OpenStudy (jennyrlz):

standard form of the equation of the parabola <--- this equation xD

OpenStudy (anonymous):

figure the rest yourself :) I can't just throw you the answer

OpenStudy (anonymous):

State the vertex and focus of the parabola having the equation (y – 3)2 = 8(x – 5). State the vertex and directrix of the parabola having the equation (x + 3)2 = –20(y – 1). The temptation is to say that the vertex is at (3, 1), but that would be wrong.

OpenStudy (jennyrlz):

u lost me with all of the "���"